Please answer all these questions
2. Maxs blood pressure reading is 145/110mmHg (1pt) a. Calculate Max s pulse pressure and mean arterial pressure. b. How would you categorize Maxs blood pressure? 3. Name two mechanisms that help blood return to the heart from the veins and explain how they work. (1pt) 4. How could extreme hypotension adversely affect a persons health? (1pt) 5. How would hypertension adversely affect a persons health? (1pt 6. How does aerobic exercise affect systolic pressure, diastolic pressure, pulse pressure, heart rate, and respiratory rate? How do those changes affect cardiac output? oxygen delivery to the blood? oxygen delivery to the tissue? (2pts)
Answer #1

2. a. Pulse pressure: 145-110 = 35

b. MAP = (SBP + 2 x DBP)/3 = (145 + 220)/3 = 365/3 = 121.7

3. The skeletal muscle pump - skeletal muscles that aid in venous return

Respiratory pump - by the pressure gradient between the intrathoracic pressure and the abdominal pressure

4. Extreme hypotension can deprive the brain and vital organs of oxygen and nutrients, leaving it in a state of shock. It can result in fainting and seizures.

5. Hypertension leads to weaker arteries, and can lead to aneurysms, coronary heart disease, stroke, heart failure among others. Most of this is caused due to higher pressure in the blood vessels, but presence of lipids clogging the vessels make the symptoms worse.